Solve equation by taking square roots.
2x^2 - 7 = 59

Answer by oberobic(2304)

2x^2 -7 = 59
.
2x^2 -66 = 0
.
x^2 -33 = 0
.
(x + sqrt(33))*(x - sqrt(33)) = 0
.
x = + or - sqrt(33)
.
Check the answer
.
x^2 = (sqrt(33)^2 = 33^(1/2)^2 = 33
2*33 = 66
66-7 = 59
.
Answer: x = + or - sqrt(33)